OpenStudy (anonymous):

What is the area of a sector with radius 10" and measure of arc equal to 45°?

OpenStudy (saifoo.khan):

45 ---- pi r^2 360

OpenStudy (saifoo.khan):

45 ---- pi 10^2 360

OpenStudy (saifoo.khan):

radius cant be 10 **degrees**

OpenStudy (saifoo.khan):

area = 39.27units dq

OpenStudy (saifoo.khan):

sq*

OpenStudy (anonymous):

45/360 = x / (pi)100 1/8 = x / 314.159... 314.159... = 8x x= 39.269"

OpenStudy (anonymous):

and because things are hard to write on computers...